A satellite circles the earth in an orbit whose radius is five times the earth's radius. The earth's mass is 5.98 1024 kg, and its radius is 6.38 106 m. What is the period of the satellite?
Answer #1

Period of satellite is given as

T = 2*pi*sqrt(r^3/(G*M)) = 2*(22/7)*sqrt((5*6.38*10^6)^3/(6.67*10^(-11)*5.98*10^24)) = 56705.7 sec